OpenStudy (ashking1):

Which ratio forms a proportion with 30/42

OpenStudy (callisto):

\[\frac{30}{42}=\frac{6\times 5}{6\times 7}=\frac{5}{ 7}\]

OpenStudy (callisto):

Ratio form: 5:7

OpenStudy (ashking1):

these are the choices 45/63 15/27 28/40 21/15

OpenStudy (callisto):

Simplify the choice like what I've done for the question and see which gives you 5/7. That's that answer you need. Can you try?

OpenStudy (ashking1):

they dont simplify to 5/7